# Mitigate Position Bias in Large Language Models via Scaling a Single Dimension

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly applied in various real-world scenarios due to their excellent generalization capabilities and robust generative abilities. However, they exhibit position bias, also known as "lost in the middle", a phenomenon that is especially pronounced in long-context scenarios, which indicates the placement of the key information in different positions of a prompt can significantly affect accuracy. This paper first explores the micro-level manifestations of position bias, concluding that attention weights are a micro-level expression of position bias. It further identifies that, in addition to position embeddings, causal attention mask also contributes to position bias by creating position-specific hidden states. Based on these insights, we propose a method to mitigate position bias by scaling this positional hidden states. Experiments on the NaturalQuestions Multi-document QA, KV retrieval, LongBench and timeline reorder tasks, using various models including RoPE models, context window-extended models, and Alibi models, demonstrate the effectiveness and generalizability of our approach. Our method can improve performance by up to 15.2% by modifying just one dimension of hidden states. Our code is available at https://aka.ms/PositionalHidden.
@@ 1 Introduction
Long-context large language models (LLMs) [1, 2, 3, 4, 5, 6] have recently garnered significant attention within the community, enabling LLMs to handle longer and more complex tasks such as long-context question-answering [7, 8] and repository-level code understanding [9]. However, recent researches [8, 10, 11, 12, 13], indicates that these long-context LLMs struggle to effectively and consistently utilize all the information provided in the context, exhibiting a position bias known as "lost in the middle", which means LLMs tend to ignore information in the middle of the prompt, even though they can utilize the information at the beginning and end of the prompts well. This issue occurs in nearly all LLMs [10, 14, 15], whether they are decoder-only models or encoder-decoder models, powerful models or small LLMs. For example, for the GPT-3.5-Turbo model in the NaturalQuestion multi-document QA task, the performance difference between ground-truth information placed in the middle of the prompt versus at the ends is 22 points with 2.3k tokens prompt [10]. This significantly impacts the practical application of LLMs in real-world scenarios. Studies [16, 17] show that this position bias becomes more severe as the context length increases, hindering the practical application of long-context LLMs.
@@
Previous works have analyzed this issue from the perspectives of data distribution [14, 18, 19] and position embeddings [15, 20]. For example, FILM [19] addresses position bias by constructing data with key information distributed in various positions for supervised fine-tuning (SFT). Ms-PoE [15] mitigates position bias by interpolating RoPE [21] using head-wise scaling factors. However, these methods require additional overhead for training or online estimation of scaling coefficients and are currently applicable to only a few models, limiting their generalizability.
@@
To fundamentally understand and alleviate position bias in LLMs, we first explored the micro-level manifestation of position bias in LLMs and observed patterns in the attention weights consistent with position bias. Next, we investigated the underlying causes of attention weight-induced position bias. By respectively modifying position embedding and causal mask, we found that, in addition to position embedding, the causal mask also significantly affects position bias. Further analysis revealed that the causal mask introduces "positional hidden states", which are positively correlated with absolute positions, thereby conveying positional information to LLMs. These positional hidden states appear regardless of what position encoding method is used, including RoPE [21], Alibi [22], and even NoPE [23].
@@
Based on the above findings, we propose a position bias mitigation method named "scale positional hidden states". Specifically, we first design a prior-based searching algorithm that quickly identifies which dimensions of hidden states within the model are positional hidden states, using monotonicity, smoothness, and loss on validation sets as indicators. Next, we design an attention modification algorithm that only let the scaled hidden states influence the attention of the last token of the prompt, efficiently implemented using FlashAttention [24].
@@
Extensive experiments on various models, including LLaMA-2 [25], Vicuna [26], Mistral [27], Gemma [28], Qwen [29], and MPT [30], and across different tasks, including Multi-document QA, KV retrieval, LongBench [31] benchmark, and the timeline reorder task [11], demonstrate that our method effectively mitigates position bias by modifying only one dimension of the hidden states of the model, achieving improvements of up to 15.2%. Our method is compatible with various position embeddings, including RoPE [21] and Alibi [22], and shows good generalization.
@@
2 Beyond Position Embeddings: Causal Masks Also Contribute to Position Bias in LLMs
This section identifies patterns in attention weights that closely correspond to position bias. Additionally, we discover that, apart from position embeddings, position information in the LLMs can also be generated by the causal mask, which tends to accumulate in a few specific hidden states channels and bears significant responsibility for the emergence of position bias.
@@ 2.1 Microscopic Manifestations of Position Bias in Transformers: Attention Weight Patterns
where $\bm{h}$ is the hidden states, and $\bm{h}(n)$ is the hidden state of the n-th token. $W^{Q},W^{K}$ are the weights of the linear layers, $\mathcal{P}$ is the position encoding function like RoPE [21], $d$ is the dimensionality of query and key states, and $n$ and $m$ are the positional order information. Mask is the causal mask.
To explore the micro-level manifestations of position bias in Transformers, we analyzed the attention weights for sentences containing key information, using a KV retrieval task, which requires the model to retrieval the ground-truth value of the given key from a list containing 50 Key-Value pairs (see Appendix D for details). As shown in Figures 1, in deep layers the model exhibits retrieval-like behavior, focusing on ground-truth information, forming a diagonal pattern observed in Figure 1b. While in other shallow layers, it always focus most attention on the start or end of the prompt, wherever the key information is located, exhibiting vertical lines patterns, as shown in Figure 1a.
In these layers exhibiting retrieval-like behavior, it can be observed that the attention weights for key information (Gold KV) exhibit patterns similar to position bias: when key information is located at the start or end of the prompt, the attention weights focused on it are relatively higher, while in the middle, they are significantly lower. Moreover, we extract the attention to key information (average of layers 15~25) with different context length in Figure 1c, where as the context length grows, the attenuation of attention weights with respect to position becomes more pronounced, reaching almost zero at the middle. More details about this are in Appendix I and D.
Furthermore, in Appendix E, we found artificially adjusting the attention weights to the key information can directly improve the corresponding accuracy. Thus, we claim that position bias is to a large extent caused by the attention weights patterns at the micro level.
@@
Figure 1: Attention distribution of the ground-truth KV pair to each KV pair across different positions on the KV retrieval task [10] using Mistral-7B [27]. (a) and (b) show the results averaged across all heads of the layer. (c) shows the attention of the ground-truth KV to the ground-truth KV (i.e., diagonal lines from (b)) across different context lengths.
@@ 2.2 Causal Mask Also Contributes to Position Bias
Figure 2: Performance of different methods with the ground-truth KV at different positions in the KV retrieval task [10] using Mistral-7B [27].
Based on Equ.(1), position embedding $\mathcal{P}$ allows LLMs to acquire postional information. However, recent works [23, 32, 33] indicate that, besides position embeddings, the causal mask can also introduce positional information.
Thereby, in this section, we aim to determine whether these two factors affect position bias through modifying different properties of the ground-truth KV pair. Specifically, we introduce the following three baselines: (1) Crop Mask, which modifies the casual mask so that the ground-truth KV pair only sees itself but not the previous tokens. (2) PE to Beginning, which reduces the position IDs of the ground-truth KV pair to be the same as the first KV pair. (3) PE to End, which increases the position IDs of the ground-truth KV pair to be the same as the last KV pair. More details are provided in Appendix F.
As shown in Figure 2, the original results exhibit a "lost in the middle" pattern not only in accuracy but also in attention weight. Secondly, PE to end has a certain degree of help, but can hardly allow the model’s performance to match the accuracy when the ground-truth KV pair is positioned at the start or end of the prompt. Furthermore, PE to Beginning results in a noticeable performance drop as well as attention weight reduction when the gold KV is close to the end. In contrast, modifying the casual mask effectively enhances attention, especially to the latter KVs, and let the performance at the middle be improved to almost on par with the beginning. Based on the above observations, we can conclude that besides position embedding, the casual mask is also an important factor affecting position bias as well as corresponding attention weights. Moreover, solely modifying the position embedding hardly alleviates position bias completely.
@@ 2.3 Casual Mask Stores Position Information in Specific Hidden states Channels
To further analyze how positional information is transmitted in transformers, we define a special type of hidden state that directly reflects absolute positional information with high correlations to position IDs, called positional hidden states, as defined in Definition 2.1. We employ monotonicity rather than correlation as the primary property of positional hidden states, as correlation does not account for the sequential nature of positions. As shown in Figure 3, our experiments reveal that causal LLMs consistently possess such hidden states across most layers (details in Appendix J), even though these models do not have explicit absolute position embeddings, which means the causal mask is a very possible factor that provides absolute positional information. We indeed prove it has major influence on positional hidden states through perturbation experiments on the causal mask and position embedding in Appendix F. So far, combining the conclusion from section 2.2, we can conclude that the causal mask stores positional information in some special hidden states, and further affects attention weights to cause position bias.
@@ 3.1 Problem Formulation
Given a pre-trained LLM $\boldsymbol{\theta}$ and a general dataset $\{\bm{x},\bm{y}\}$, our objective is to find the optimal positional hidden states $h_{t}$ and the corresponding scaling factor $s$ to maximally reduce position bias, which can be formulated as follows:
$\mathop{\arg\min}_{h_{t}\in\mathcal{H},s<1}\mathbb{E}\left[\sum_{i=1}^{|\bm{P}|}\mathcal{L}\left(\bm{x},\bm{y},\bm{p}_{i};F(\boldsymbol{\theta},h_{t},s)\right)\right]$ (2)
where $\bm{P}$ represents the set of different positions of the ground-truth information within the prompt $\bm{x}$, $F(\boldsymbol{\theta},h_{t},s)$ denotes the operation of scaling the LLM $\boldsymbol{\theta}$ on the $t$-th dimension of its hidden states by the scaling factor $s$, and $\mathcal{L}$ denotes the loss for general downstream tasks of the modified model.
@@ 3.2 Identifying Positional Hidden States
To efficiently search for the positional hidden states from the LLMs’ hidden states set, we leverage the characteristics of positional hidden states defined in Section 2.3 and propose a prior-based positional hidden search algorithm. As shown in Algorithm 1, the search process consists of the following two steps: 1) Identify the top-$k$ dimensions $\rho$ in the hidden states that are monotonic in more than $\varepsilon$ layers and are as smooth as possible. Here $c_{t}$ is the number of layers where $h_{t}(p)$ is monotonic, and $g_{t}$ is the smooth score of $h_{t}(p)$. Equ.(3) is the smoothness formula. 2) Use a small validation dataset $\mathcal{D}_{\text{val}}=\{\bm{x},\bm{y}\}$ to evaluate the impact of scaling these positional hidden states respectively and select the positional hidden states $h_{\overline{t}}$ that can lead to the minimal loss $\mathcal{L}_{\overline{t}}$.
@@ 3.3 Scaling the Positional Hidden States
To minimize the impact of this modification on the semantics of LLMs, we propose a method scaling the positional hidden states only affecting the last token as shown in Figure 4. Specifically, for the tokens preceding the last token, the attention calculation remains the same as the original. For the last token’s attention computation of a sequence of length $l$, we obtain the modified query state $\bm{\overline{q}}_{l}$ (of the $l$-th token, i.e. the last token) and key states $\bm{\overline{K}}$ (of all the tokens) by scaling the positional hidden states. That is,
@@
where ${\bm{z}}$ is the attention output. We use FlashAttention [24] to implement our method with minimal overhead. After obtaining the combined attention weights, the remaining computations are same with the original. As shown in Appendix C.4, our method only causes a slight increase in latency.
@@ 4.1 Setup
We apply our method to a wide range of state-of-the-art open-source LLMs, including: 1) RoPE [34] models: LLaMA-2 (7B, 13B) [25], Mistral-7B [27], Gemma-7B [28], Qwen1.5-7B [29]; 2) Context window extended models: Vicuna (7B, 13B) [26]; 3) Alibi [22] models: MPT-30B [30]. All the models we use are instruction-tuned versions.
And we evaluate the performance across three aspects: 1) Position-bias-related tests on NaturalQuestion multi-document QA [10] and KV retrieval [10] with ground-truth at different positions in the prompt. The NaturalQuestion task includes 20 documents with a prompt length of about 2.3k tokens, while the KV retrieval task includes 140 KV pairs with an average length of about 10k tokens. 2) General long-context benchmark on LongBench [31], including multi-document QA, single-document QA, summarization, few-shot learning, synthetic tasks, and code completion, totaling 16 tasks with an average length of 37k tokens. 3) Position-sensitive tasks on timeline reordering in LooGLE [11], with an average length of 10k tokens. For prompts that exceed the context windows of LLMs, we follow LongBench’s approach by truncating from the middle and retaining the head and tail of the prompt to fit within the context windows. We use the provided metrics and scripts from the following benchmarks for evaluation.
@@
In this paper, we implement our approach using PyTorch, HuggingFace Transformers, and FlashAttention [24] in an A100 GPU. To ensure stable and reproducible results, we use greedy decoding in all experiments. For the search part, we set the top-$k$ size of positional hidden states to 10 and $\varepsilon$ to $L/4$, where $L$ is the number of layers. The validation set is a synthetic KV retrieval dataset consisting of 100 examples, which do not overlap with the test set. The search process takes approximately 10 minutes. For the scaling part, we only modify the intermediate layers of the model to minimize the negative impact on performance. The details of the scaling dimensions, layer ranges, and factors are shown in Table 4. More details are provided in Appendix C.
@@
We include two training-free positional bias mitigation approaches as our baselines: (i) Original, the original results of LLMs with the ground-truth at different positions in the prompt. (ii) w/ Ms-PoE [15], uses a head-aware position embedding scaling method to mitigate position bias. We follow the paper’s settings and apply scaling coefficients of 1.2 to 1.8 from the $3$-rd layer.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly applied in various real-world scenarios due to their excellent generalization capabilities and robust generative abilities. However, they exhibit position bias, also known as "lost in the middle", a phenomenon that is especially pronounced in long-context scenarios, which indicates the placement of the key information in different positions of a prompt can significantly affect accuracy. This paper first explores the micro-level manifestations of position bias, concluding that attention weights are a micro-level expression of position bias. It further identifies that, in addition to position embeddings, causal attention mask also contributes to position bias by creating position-specific hidden states. Based on these insights, we propose a method to mitigate position bias by scaling this positional hidden states. Experiments on the NaturalQuestions Multi-document QA, KV retrieval, LongBench and timeline reorder tasks, using various models including RoPE models, context window-extended models, and Alibi models, demonstrate the effectiveness and generalizability of our approach. Our method can improve performance by up to 15.2% by modifying just one dimension of hidden states.
@@ 1 Introduction
Long-context large language models (LLMs) (Gradient, 2024; Reid et al., 2024; Liu et al., 2024a; Young et al., 2024; Abdin et al., 2024; DeepSeek-AI, 2024) have recently garnered significant attention within the community, enabling LLMs to handle longer and more complex tasks such as long-context question-answering (Caciularu et al., 2023; Li et al., 2024b) and repository-level code understanding (Bairi et al., 2024). However, recent researches (Li et al., 2024b; Liu et al., 2024c; Li et al., 2024a; Shi et al., 2023; Tang et al., 2023), indicates that these long-context LLMs struggle to effectively and consistently utilize all the information provided in the context, exhibiting a position bias known as "lost in the middle", which means LLMs tend to ignore information in the middle of the prompt, even though they can utilize the information at the beginning and end of the prompts well. This issue occurs in nearly all LLMs (Liu et al., 2024c; Junqing et al., 2023; Zhang et al., 2024), whether they are decoder-only models or encoder-decoder models, powerful models or small LLMs. For example, for the GPT-3.5-Turbo model in the NaturalQuestion multi-document QA task, the performance difference between ground-truth information placed in the middle of the prompt versus at the ends is 22 points with 2.3k tokens prompt (Liu et al., 2024c). This significantly impacts the practical application of LLMs in real-world scenarios. Studies (Kamradt, 2023; Zhao et al., 2024) show that this position bias becomes more severe as the context length increases, hindering the practical application of long-context LLMs.
@@
Previous works have analyzed this issue from the perspectives of data distribution (Junqing et al., 2023; Yu, 2023; An et al., 2024) and position embeddings (Zhang et al., 2024; Chen et al., 2024). For example, FILM (An et al., 2024) addresses position bias by constructing data with key information distributed in various positions for supervised fine-tuning (SFT). Ms-PoE (Zhang et al., 2024) mitigates position bias by interpolating RoPE (Su et al., 2024) using head-wise scaling factors. However, these methods require additional overhead for training or online estimation of scaling coefficients and are currently applicable to only a few models, limiting their generalizability.
@@
To fundamentally understand and alleviate position bias in LLMs, we first explored the micro-level manifestation of position bias in LLMs and observed patterns in the attention weights consistent with position bias. Next, we investigated the underlying causes of attention weight-induced position bias. By respectively modifying position embedding and causal mask, we found that, in addition to position embedding, the causal mask also significantly affects position bias. Further analysis revealed that the causal mask introduces "positional hidden states", which are positively correlated with absolute positions, thereby conveying positional information to LLMs. These positional hidden states appear regardless of what position encoding method is used, including RoPE (Su et al., 2024), Alibi (Press et al., 2022), and even NoPE (Haviv et al., 2022).
@@
Based on the above findings, we propose a position bias mitigation method named "scale positional hidden states". Specifically, we first design a prior-based searching algorithm that quickly identifies which dimensions of hidden states within the model are positional hidden states, using monotonicity, smoothness, and loss on validation sets as indicators. Next, we design an attention modification algorithm that only let the scaled hidden states influence the attention of the last token of the prompt, efficiently implemented using FlashAttention (Dao, 2024).
@@
Extensive experiments on various models, including LLaMA-2 (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023), Gemma (Team et al., 2024), Qwen (Bai et al., 2023), and MPT (Team, 2023), and across different tasks, including Multi-document QA, KV retrieval, LongBench (Bai et al., 2024) benchmark, and the timeline reorder task (Li et al., 2024a), demonstrate that our method effectively mitigates position bias by modifying only one dimension of the hidden states of the model, achieving improvements of up to 15.2%. Our method is compatible with various position embeddings, including RoPE (Su et al., 2024) and Alibi (Press et al., 2022), and shows good generalization.
@@
2 Beyond Position Embeddings: Positional Information can be seen in hidden states
In this section, we first identifies patterns in attention weights that closely correspond to position bias. Then, we discover that, apart from position embeddings, position information in the LLMs can also be generated by the causal mask, which tends to accumulate in a few specific hidden states channels and bears significant responsibility for the emergence of position bias.
@@ 2.1 Microscopic Manifestations of Position Bias in Transformers: Attention Weight Patterns
where $\bm{h}$ is the hidden states, and $\bm{h}(n)$ is the hidden state of the n-th token. $W^{Q},W^{K}$ are the weights of the linear layers, $\mathcal{P}$ is the position encoding function like RoPE (Su et al., 2024), $d$ is the dimensionality of query and key states, and $n$ and $m$ are the positional order information. Mask is the causal mask.
To explore the micro-level manifestations of position bias in Transformers, we analyzed the attention weights for sentences containing key information, using a KV retrieval task, which requires the model to retrieval the ground-truth value of the given key from a list containing 50 Key-Value pairs (see Appendix B for details). As shown in Figures 1, in deep layers the model exhibits retrieval-like behavior, focusing on ground-truth information, forming a diagonal pattern observed in Figure 1b. While in other shallow layers, it always focus most attention on the start or end of the prompt, wherever the key information is located, exhibiting vertical lines patterns, as shown in Figure 1a.
In these layers exhibiting retrieval-like behavior, it can be observed that the attention weights for key information (Gold KV) exhibit patterns similar to position bias: when key information is located at the start or end of the prompt, the attention weights focused on it are relatively higher, while in the middle, they are significantly lower. Moreover, we extract the attention to key information (average of layers 15~25) with different context length in Figure 1c, where as the context length grows, the attenuation of attention weights with respect to position becomes more pronounced, reaching almost zero at the middle. More details about this are in Appendix E and B.
Furthermore, in Appendix 4.3, we found artificially adjusting the attention weights to the key information can directly improve the corresponding accuracy. Thus, we claim that position bias is to a large extent caused by the attention weights patterns at the micro level.
@@
Figure 1: Attention distribution of the ground-truth KV pair to each KV pair across different positions on the KV retrieval task (Liu et al., 2024c) using Mistral-7B (Jiang et al., 2023). (a) and (b) show the results averaged across all heads of the layer. (c) shows the attention of the ground-truth KV to the ground-truth KV (i.e., diagonal lines from (b)) across different context lengths.
@@ 2.2 Causal Mask Also Contributes to Position Bias
Figure 2: Performance of different methods with the ground-truth KV at different positions in the KV retrieval task (Liu et al., 2024c) using Mistral-7B (Jiang et al., 2023).
Based on Equ.(1), position embedding $\mathcal{P}$ allows LLMs to acquire postional information. However, recent works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) indicate that, besides position embeddings, the causal mask can also introduce positional information.
In this section, we aim to determine whether these two factors affect position bias by modifying different properties of the ground-truth KV pair. We introduce three baselines: (1) Crop Mask, which alters the causal mask so the ground-truth KV pair sees only itself, not previous tokens; (2) PE to Beginning, which assigns the position IDs of the ground-truth KV pair to match the first KV pair; (3) PE to End, which assigns the position IDs to match the last KV pair, details in Appendix C.
As shown in Figure 2, the original results exhibit a ”lost in the middle" pattern not only in accuracy but also in attention weight. Secondly, PE to end has a certain degree of help, but can hardly allow the model’s performance to match the accuracy when the ground-truth KV pair is positioned at the start or end of the prompt. Furthermore, PE to Beginning results in a noticeable performance drop as well as attention weight reduction when the gold KV is close to the end. In contrast, modifying the casual mask effectively enhances attention, especially to the latter KVs, and let the performance at the middle be improved to almost on par with the beginning. Based on the above observations, we can conclude that besides position embedding, the casual mask is also an important factor affecting position bias as well as corresponding attention weights. Moreover, solely modifying the position embedding hardly alleviates position bias completely.
@@ 2.3 Casual Mask Stores Position Information in Specific Hidden states Channels
Previous works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) have found that the positional information generated by causal mask is implicitly stored in hidden states. However, in fact, we find it can be observed explicitly, from “positional hidden states”.
To further analyze how positional information is transmitted in transformers, we define a special type of hidden state that directly reflects absolute positional information with high correlations to position IDs, called positional hidden states, as defined in Definition 2.1. We employ monotonicity rather than correlation as the primary property of positional hidden states, as correlation does not account for the sequential nature of positions. As shown in Figure 3, our experiments reveal that causal LLMs consistently possess such hidden states across most layers (details in Appendix F), even though these models do not have explicit absolute position embeddings, which means the causal mask is a very possible factor that provides absolute positional information. To demonstrate that these position hidden states are formed under the influence of the causal mask rather than the position embeddings, we conduct perturbation experiments on the causal mask and position embedding, as shown in Appendix C.
Based on the findings from Section 2.2, we conclude that the causal mask encodes positional information in certain hidden states, which subsequently influences attention weights and introduces position bias.
@@ 3.1 Problem Formulation
Given a pre-trained LLM $\bm{\theta}$ and a general dataset $\{\bm{x},\bm{y}\}$, our objective is to find the optimal positional hidden states $h_{t}$ and the corresponding scaling factor $s$ to maximally reduce position bias, which can be formulated as follows:
$\mathop{\arg\min}_{h_{t}\in\mathcal{H},s<1}\mathbb{E}\left[\sum_{i=1}^{|\bm{P}|}\mathcal{L}\left(\bm{x},\bm{y},\bm{p}_{i};F(\bm{\theta},h_{t},s)\right)\right]$ (2)
where $\bm{P}$ represents the set of different positions of the ground-truth information within the prompt $\bm{x}$, $F(\bm{\theta},h_{t},s)$ denotes the operation of scaling the LLM $\bm{\theta}$ on the $t$-th dimension of its hidden states by the scaling factor $s$, and $\mathcal{L}$ denotes the loss for general downstream tasks of the modified model.
@@ 3.2 Identifying Positional Hidden States
To efficiently search for the positional hidden states from the LLMs’ hidden states set, we leverage the characteristics of positional hidden states defined in Section 2.3 and propose a heuristic positional hidden search algorithm. As shown in Algorithm 1, the search process consists of the following two steps: 1) Identify the top-$k$ dimensions $\rho$ in the hidden states that are monotonic in more than $\varepsilon$ layers and are as smooth as possible. Here $c_{t}$ is the number of layers where $h_{t}(p)$ is monotonic, and $g_{t}$ is the smooth score of $h_{t}(p)$. Equ.(3) is the smoothness formula. 2) Use a small validation dataset (or called calibration dataset) $\mathcal{D}_{\text{val}}=\{\bm{x},\bm{y}\}$ to evaluate the impact of scaling these positional hidden states respectively and select the positional hidden states $h_{\overline{t}}$ that can lead to the minimal loss $\mathcal{L}_{\overline{t}}$.
@@ 3.3 Scaling the Positional Hidden States
To minimize the impact of this modification on the semantics of LLMs, we propose a method scaling the positional hidden states only affecting the last token, as shown in Figure 4. Specifically, for the tokens preceding the last token, the attention calculation remains the same as the original. For the last token’s attention computation of a sequence of length $l$, we obtain the modified query state $\bm{\overline{q}}_{l}$ (of the $l$-th token, i.e. the last token) and key states $\bm{\overline{K}}$ (of all the tokens) by scaling the positional hidden states. That is,
@@
where ${\bm{z}}$ is the attention output. We implement our method using FlashAttention (Dao, 2024) with minimal overhead. After calculating the combined attention weights, the remaining computations remain the same as in the original method. As shown in Appendix A.4, our approach results in only a slight increase in latency.
@@ 4.1 Setup
We apply our method to a wide range of state-of-the-art open-source LLMs, including: 1) RoPE (Chen et al., 2023) models: LLaMA-2 (7B, 13B) (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma-7B (Team et al., 2024), Qwen1.5-7B (Bai et al., 2023); 2) Context window extended models: Vicuna (7B, 13B) (Chiang et al., 2023); 3) Alibi (Press et al., 2022) models: MPT-30B (Team, 2023). All the models we use are instruction-tuned versions.
And we evaluate the performance across three aspects: 1) Position-bias-related tests on NaturalQuestion multi-document QA (Liu et al., 2024c) and KV retrieval (Liu et al., 2024c) with ground-truth at different positions in the prompt. The NaturalQuestion task includes 20 documents with a prompt length of about 2.3k tokens, while the KV retrieval task includes 140 KV pairs with an average length of about 10k tokens. 2) General long-context benchmark on LongBench (Bai et al., 2024), including multi-document QA, single-document QA, summarization, few-shot learning, synthetic tasks, and code completion, totaling 16 tasks with an average length of 37k tokens. 3) Position-sensitive tasks on timeline reordering in LooGLE (Li et al., 2024a), with an average length of 10k tokens. For prompts that exceed the context windows of LLMs, we follow LongBench’s approach by truncating from the middle and retaining the head and tail of the prompt to fit within the context windows. We use the provided metrics and scripts from the following benchmarks for evaluation.
@@
In this paper, we implement our approach using PyTorch, HuggingFace Transformers, and FlashAttention (Dao, 2024) in an A100 GPU. To ensure stable and reproducible results, we use greedy decoding in all experiments. For the search part, we set the top-$k$ size of positional hidden states to 10 and $\varepsilon$ to $L/4$, where $L$ is the number of layers. The validation set is a synthetic KV retrieval dataset consisting of 100 examples, which do not overlap with the test set. The search process takes approximately 10 minutes. For the scaling part, we only modify the intermediate layers of the model to minimize the negative impact on performance. The details of the scaling dimensions, layer ranges, and factors are shown in Table 5. More details are provided in Appendix A.
@@
We include two training-free positional bias mitigation methods as baselines: (i) Original, the unmodified LLM results with the ground truth at different positions in the prompt. (ii) w/ Ms-PoE (Zhang et al., 2024), a head-aware position embedding scaling method to mitigate position bias. Following the original settings, we apply scaling coefficients of 1.2 to 1.8 starting from the 3rd layer.